OpenStudy (anonymous):

Help!!!! a missile is launched from the ground. its height, h(x), can be represented by a quadratic function in terms of time, x, in seconds. after 1 second, the missile is 103 feet in the air; after 2 seconds, it is 240 feet in the air. find the height, in feet, of the missile after 11 seconds in the air.

OpenStudy (anonymous):

@hotdancingbabe14

OpenStudy (xapproachesinfinity):

eh let's try tackle this! can we? @lexikaylynn

OpenStudy (xapproachesinfinity):

we recall that the quadratic function is written h(x)=ax^2+bx+c in general

OpenStudy (anonymous):

ya sorry i didnt know you replied to this lol

OpenStudy (anonymous):

@iGreen or @micahm can you help me?

OpenStudy (micahm):

name of lesson

OpenStudy (anonymous):

solving problems with quadratic functions

OpenStudy (micahm):

i don't know this type of math sorry

OpenStudy (anonymous):

:( dam @iGreen do you know this?

OpenStudy (xapproachesinfinity):

I'm thinking that at t=0 the height is also zero (they said the ground level) then we conclude that c=0 so we have h(x)=ax^2+bx if we put what we now already h(1)=a+b=103 h(2)=4a+2b=240 we find a and b then after that we can find the height for any time

OpenStudy (anonymous):

i dont get how that makes scene?? :o

OpenStudy (xapproachesinfinity):

hehe first of all you agree that h(x)=ax^2+bx+c